The product of 7 and a number is 2 less than the number. What is the "number"? I know the answer is -1/3 my problem is how to derive that answer? Thank you very much for your help!
Answer by mananth(16949)

let the number be x
product of 7 & x= 7x
7x=x-2
add -x to both sides
7x-x=x-x-2
6x=-2
divide by 6
6x/6= -2/6
x= -1/3